Red, Inc., Yellow Corp., and Blue Company each will pay a dividend of $3.05 next year. The growth rate in dividends for all three companies is 6 percent. The required return for each company’s stock is 9 percent, 12 percent, and 15 percent, respectively. What is the stock price for each company? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
1
Expert's answer
2020-05-07T08:35:41-0400

according to the model of Gordon:

"P=\frac{d}{r-g}"


1."P=\frac{3.05}{9-6}=101.7"

2."P=\frac{3.05}{12-6}=50.8"

3."P=\frac{3.05}{15-6}=33.9"